# Barron Space Representations for Elliptic PDEs with Homogeneous Boundary Conditions

### v1

@@ 2.3 Main Results
Assumption 2.11 (Coefficients for Dirichlet Problem).
@@
Assumption 2.12 (Coefficients for Neumann Problem).
@@
Theorem 2.13 (Main Result 1, Formal Version).
@@
    Let $u_{D}^{*}\in H_{0}^{1}(\Omega)$ denote the weak solution to the Dirichlet problem (1.3). Suppose Assumptions 2.1 and 2.11 hold. Then for any $\varepsilon\in(0,2/\lambda_{\min})$, there exists a two-layer neural network of the form
    $N_{k}(x)\coloneqq\frac{1}{k}\sum_{i=1}^{k}a_{i}\cos\left(w_{i}^{\top}x+b_{i}\right),\quad k\leq\left\lceil\frac{\alpha_{\ast}^{2}\ell_{f,D}^{2}\left(p_{D}(d)^{T}-1\right)^{2}}{\varepsilon^{2}\left(p_{D}(d)-1\right)^{2}}\right\rceil,$
@@
    Let $u_{N}^{*}\in H^{1}(\Omega)$ denote the weak solution to the Neumann problem (1.4). Suppose Assumptions 2.1 and 2.12 hold. Then for any $\varepsilon\in(0,2/\lambda_{\min})$, there exists a two-layer neural network of the form
    $N_{k}(x)\coloneqq\frac{1}{k}\sum_{i=1}^{k}a_{i}\cos\left(w_{i}^{\top}x+b_{i}\right),\quad k\leq\left\lceil\frac{\alpha_{\ast}^{2}\ell_{f,N}^{2}\left(p_{N}(d)^{T}-1\right)^{2}}{\varepsilon^{2}\left(p_{N}(d)-1\right)^{2}}\right\rceil,$
@@
In other words, in both cases, the solution can be approximated in the $H^{1}(\Omega)$-norm by a two-layer neural network with cosine activation using at most $\mathcal{O}(d^{C|\log\varepsilon|})$ neurons, where the constant $C>0$ depends only on the coefficients of the PDE.
@@
Theorem 2.17 (Main Result 2, Formal Version).
@@
    Let $u_{D}^{*}\in H_{0}^{1}(\Omega)$ denote the weak solution to the Dirichlet problem (1.3). Suppose Assumptions 2.1 and 2.11 hold. Then for any $\varepsilon\in(0,2/\lambda_{\min})$, there exists a two-layer neural network of the form
@@
    Let $u_{N}^{*}\in H^{1}(\Omega)$ denote the weak solution to the Neumann problem (1.4). Suppose Assumptions 2.1 and 2.12 hold. Then for any $\varepsilon\in(0,2/\lambda_{\min})$, there exists a two-layer neural network of the form
@@
Here, $q(d)=256d^{2}+128d+4$, and all parameters are as specified in Remark 2.14. Thus, in both cases, the solution can be approximated in the $H^{1}(\Omega)$-norm by a two-layer neural network with $\operatorname{ReLU}$ activation using at most $\mathcal{O}(d^{C|\log\varepsilon|})$ neurons, where the constant $C>0$ depends only on the coefficients of the PDE.
Remark 2.18.
A similar approximation result for $\operatorname{ReLU}$ activation was also given in [LLW21]. However, our result is stronger and more general in the following aspects:
@@ 3 Proof of the Main Results
The objective of this section is to prove Theorems 2.13 and 2.17, following the three-step outline presented in Section 1.1. Our argument builds upon the methodology developed in [CLL21, LLW21, MLLR23, CLLZ24].
@@ 3.1.1 The Construction
    $u_{D}^{*}=\mathop{\operatorname{arg\,min}}_{u\in H_{0}^{1}(\Omega)}\mathcal{E}_{D}(u),\quad\text{where}\quad\mathcal{E}_{D}(u)\coloneqq\int_{\Omega}\frac{1}{2}\nabla u\cdot A\nabla u+\frac{1}{2}cu^{2}-fu\,\mathrm{d}x.$ (3.1)
@@
    $u_{N}^{*}=\mathop{\operatorname{arg\,min}}_{u\in H^{1}(\Omega)}\mathcal{E}_{N}(u),\quad\text{where}\quad\mathcal{E}_{N}(u)\coloneqq\int_{\Omega}\frac{1}{2}\nabla u\cdot A\nabla u+\frac{1}{2}cu^{2}-fu\,\mathrm{d}x.$ (3.2)
@@
The classical gradient descent method iteratively updates an initial guess by moving in the direction of the negative gradient of the objective function, aiming to converge to a minimizer. In the context of the PDEs we consider, Proposition 3.1 asserts that the weak solution is the unique minimizer of the associated energy functional. This naturally motivates the application of a gradient descent-type approach to solve the PDE. However, in contrast to the finite-dimensional setting, where the gradient is readily defined via classical derivatives, the notion of a gradient is not immediately available for energy functionals defined on infinite-dimensional spaces.
To overcome this difficulty, we recall an alternative way to characterize gradients. For a real-valued continuously differentiable function $g\colon\Omega\to\mathbb{R}$, its gradient $\nabla g(x)$ is the unique function $\nabla g(x)\colon\Omega\to\mathbb{R}^{d}$ satisfying
[…]
where $(\cdot,\cdot)_{\mathbb{R}^{d}}$ denotes the standard inner product in $\mathbb{R}^{d}$.
In our setting, taking the energy functional $\mathcal{E}_{D}$ for the Dirichlet problem as an example, we note that the inner product, subtraction, and scalar multiplication in (3.3) all have natural counterparts in the Hilbert space $H_{0}^{1}(\Omega)$. This motivates us to define the gradient of $\mathcal{E}_{D}$ as the map
@@
$(D_{H_{0}^{1}}\mathcal{E}_{D}(u),v)_{H^{1}(\Omega)}=\lim_{t\to 0}\frac{\mathcal{E}_{D}(u+tv)-\mathcal{E}_{D}(u)}{t}\quad\text{for all }u,v\in H_{0}^{1}(\Omega).$ (3.4)
Substituting the expression of $\mathcal{E}_{D}(u)$ from (3.1) into (3.4), we obtain
$\left(D_{H_{0}^{1}}\mathcal{E}_{D}(u),v\right)_{H^{1}(\Omega)}=\int_{\Omega}\nabla u\cdot A\nabla v+cuv-fv\,\mathrm{d}x,\quad\text{for all }u,v\in H_{0}^{1}(\Omega).$
Here we use the inner product on $H_{0}^{1}(\Omega)$ inherited from the standard inner product on $H^{1}(\Omega)$. The uniqueness of $D_{H_{0}^{1}}\mathcal{E}_{D}$ follows from the Riesz representation theorem in the Hilbert space $H_{0}^{1}(\Omega)$, applied to the linear functional
$H_{0}^{1}(\Omega)\longrightarrow\mathbb{R},\quad v\longmapsto\int_{\Omega}\nabla u\cdot A\nabla v+cuv-fv\,\mathrm{d}x,$
for each $u\in H_{0}^{1}(\Omega)$.
A similar argument applies to the energy functional $\mathcal{E}_{N}$ for the Neumann problem. We summarize these results in the following definition.
Definition 3.2.
  1. (1)
    There exists a unique operator $D_{H_{0}^{1}}\mathcal{E}_{D}\colon H_{0}^{1}(\Omega)\to H_{0}^{1}(\Omega)$ such that
    $\left(D_{H_{0}^{1}}\mathcal{E}_{D}(u),v\right)_{H^{1}(\Omega)}=\int_{\Omega}\nabla u\cdot A\nabla v+cuv-fv\,\mathrm{d}x,\quad\text{for all }u,v\in H_{0}^{1}(\Omega).$ (3.5)
    We call $D_{H_{0}^{1}}\mathcal{E}_{D}$ the Sobolev gradient of $\mathcal{E}_{D}$ with respect to $H_{0}^{1}(\Omega)$.
  2. (2)
    There exists a unique operator $D_{H^{1}}\mathcal{E}_{N}\colon H^{1}(\Omega)\to H^{1}(\Omega)$ such that
    $\left(D_{H^{1}}\mathcal{E}_{N}(u),v\right)_{H^{1}(\Omega)}=\int_{\Omega}\nabla u\cdot A\nabla v+cuv-fv\,\mathrm{d}x,\quad\text{for all }u,v\in H^{1}(\Omega).$ (3.6)
    We call $D_{H^{1}}\mathcal{E}_{N}$ the Sobolev gradient of $\mathcal{E}_{N}$ with respect to $H^{1}(\Omega)$.
Remark 3.3.
The concept of the Sobolev gradient has been widely employed in the study of PDEs, including applications to the Gross–Pitaevskii eigenvalue problem [HP20, Zha22, CLLZ24].
Analogous to the fact that the gradient of a function vanishes at its local minimum, the Sobolev gradient of an energy functional should also vanish at its minimizer. Indeed, by (2.1) and (2.2), we have
$\left(D_{H_{0}^{1}}\mathcal{E}_{D}(u^{*}_{D}),v\right)_{H^{1}(\Omega)}$ $=0\quad\text{for all }v\in H_{0}^{1}(\Omega),$
$\left(D_{H^{1}}\mathcal{E}_{N}(u^{*}_{N}),v\right)_{H^{1}(\Omega)}$ $=0\quad\text{for all }v\in H^{1}(\Omega).$
By the non-degeneracy of the inner products on $H_{0}^{1}(\Omega)$ and $H^{1}(\Omega)$, it follows that
@@ 3.1.2 Convergence Analysis
Assume Assumption 2.1 holds. The Sobolev gradient flow (3.8) for the Dirichlet problem (1.3) converges exponentially in the $H^{1}(\Omega)$-norm to the weak solution $u_{D}^{*}$. More precisely, there exists a constant $0<\beta<1$ such that
@@
Remark 3.5.
Let
and define the step size and contraction factor by
This choice of parameters is shown to be optimal for the Sobolev gradient flow (3.8) and (3.9) in the proof of Theorem 3.4 below. We therefore adopt this choice for all subsequent analysis after the proof.
The proof of Theorem 3.4 makes use of the following well-known lemma. For completeness, we include its proof in Appendix A.2.
Lemma 3.6.
Assume that the coefficient matrix $A(x)$ satisfies the condition in Assumption 2.1. Then, for all $u,v\in H^{1}(\Omega)$, we have
$\int_{\Omega}\nabla u\cdot A\nabla v\,{\,\mathrm{d}}x\leq a_{\max}\|\nabla u\|_{L^{2}(\Omega)}\|\nabla v\|_{L^{2}(\Omega)}.$
Proof of Theorem 3.4.
We provide the proof only for the Dirichlet problem (1.3), as the proof for the Neumann problem (1.4) follows analogously. Our goal is to establish that there exists a constant $0<\beta<1$ such that
@@
$\|u_{t+1}-u_{D}^{*}\|_{H^{1}(\Omega)}^{2}$ $=\|u_{t}-u_{D}^{*}-\alpha D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{H^{1}(\Omega)}^{2}$
$=\|u_{t}-u_{D}^{*}\|_{H^{1}(\Omega)}^{2}+2\alpha\left(u_{D}^{*}-u_{t},D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\right)_{H^{1}(\Omega)}+\alpha^{2}\|D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{H^{1}(\Omega)}^{2}.$ (3.11)
@@
Let $\lambda_{\max}\coloneqq\max\{a_{\max},c_{\max}\}$. Then, by combining (3.14), the definition of the Sobolev gradient in (3.5), Lemma 3.6, and the Cauchy–Schwarz inequality, we obtain
$\|D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{H^{1}(\Omega)}^{2}=\left(D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t}),D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\right)_{H^{1}(\Omega)}-\left(D_{H_{0}^{1}}\mathcal{E}_{D}(u_{D}^{*}),D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\right)_{H^{1}(\Omega)}$
@@
$\leq$ $a_{\max}\|\nabla(u_{t}-u_{D}^{*})\|_{L^{2}(\Omega)}\|\nabla D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{L^{2}(\Omega)}+c_{\max}\|u_{t}-u_{D}^{*}\|_{L^{2}(\Omega)}\|D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{L^{2}(\Omega)}$
$\leq$ $\lambda_{\max}\left(\|\nabla(u_{t}-u_{D}^{*})\|_{L^{2}(\Omega)}\|\nabla D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{L^{2}(\Omega)}+\|u_{t}-u_{D}^{*}\|_{L^{2}(\Omega)}\|D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{L^{2}(\Omega)}\right)$
$\leq$ $\lambda_{\max}\left(\|\nabla(u_{t}-u_{D}^{*})\|_{L^{2}(\Omega)}^{2}+\|u_{t}-u_{D}^{*}\|_{L^{2}(\Omega)}^{2}\right)^{1/2}\left(\|\nabla D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{L^{2}(\Omega)}^{2}+\|D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{L^{2}(\Omega)}^{2}\right)^{1/2}$
@@
[…]
Then for sufficiently small $\alpha>0$, we can ensure that $0<\beta<1$, which completes the proof.
The optimality of the step size and contraction factor in (3.10) follows directly from analyzing the quadratic function in (3.16). ∎
@@
$\lambda_{\min}\|u_{D}^{*}\|_{H^{1}(\Omega)}^{2}\leq\int_{\Omega}\nabla u_{D}^{*}\cdot A\nabla u_{D}^{*}\,\mathrm{d}x+\int_{\Omega}c|u_{D}^{*}|^{2}\,\mathrm{d}x=\int_{\Omega}fu_{D}^{*}\,\mathrm{d}x\leq\|f\|_{H^{-1}(\Omega)}\|u_{D}^{*}\|_{H^{1}(\Omega)}.$ (3.17)
@@
then the iterates $u_{T}$ produced by both schemes (3.8) and (3.9) satisfy

### v2

@@ 2.2 Main Results
Assumption 2.5 (Coefficients of Dirichlet Problem).
@@
Assumption 2.6 (Coefficients of Neumann Problem).
@@
Theorem 2.7 (Main Result 1).
@@
    Let $u_{D}^{*}\in H_{0}^{1}(\Omega)$ denote the weak solution to the Dirichlet problem (1.3). Suppose Assumptions 2.1 and 2.5 hold. Then for any $\varepsilon\in(0,2/\lambda_{\min})$, there exists a two-layer neural network of the form
    $N_{k}(x)\coloneqq\frac{1}{k}\sum_{i=1}^{k}a_{i}\cos(w_{i}^{\top}x+b_{i}),\quad k\leq\left\lceil\frac{\alpha_{\ast}^{2}\ell_{f,D}^{2}\left(p_{D}(d)^{T}-1\right)^{2}}{\varepsilon^{2}\left(p_{D}(d)-1\right)^{2}}\right\rceil,$
@@
    Let $u_{N}^{*}\in H^{1}(\Omega)$ denote the weak solution to the Neumann problem (1.4). Suppose Assumptions 2.1 and 2.6 hold. Then for any $\varepsilon\in(0,2/\lambda_{\min})$, there exists a two-layer neural network of the form
    $N_{k}(x)\coloneqq\frac{1}{k}\sum_{i=1}^{k}a_{i}\cos(w_{i}^{\top}x+b_{i}),\quad k\leq\left\lceil\frac{\alpha_{\ast}^{2}\ell_{f,N}^{2}\left(p_{N}(d)^{T}-1\right)^{2}}{\varepsilon^{2}\left(p_{N}(d)-1\right)^{2}}\right\rceil,$
@@
In other words, in both cases, the solution can be approximated in the $H^{1}(\Omega)$-norm by a two-layer neural network with cosine activation using at most $\mathcal{O}(d^{C|\log\varepsilon|})$ neurons, where the constant $C>0$ depends only on the Barron norms of PDE coefficients.
@@
Theorem 2.9 (Main Result 2).
@@
    Let $u_{D}^{*}\in H_{0}^{1}(\Omega)$ denote the weak solution to the Dirichlet problem (1.3). Suppose Assumptions 2.1 and 2.5 hold. Then for any $\varepsilon\in(0,2/\lambda_{\min})$, there exists a two-layer neural network of the form
@@
    Let $u_{N}^{*}\in H^{1}(\Omega)$ denote the weak solution to the Neumann problem (1.4). Suppose Assumptions 2.1 and 2.6 hold. Then for any $\varepsilon\in(0,2/\lambda_{\min})$, there exists a two-layer neural network of the form
@@
Here, $q(d)=256d^{2}+128d+4$, and all parameters are as specified in Remark 2.8. Thus, in both cases, the solution can be approximated in the $H^{1}(\Omega)$-norm by a two-layer neural network with $\operatorname{ReLU}$ activation using at most $\mathcal{O}(d^{C|\log\varepsilon|})$ neurons, where the constant $C>0$ depends only on the Barron norms of PDE coefficients.
A similar approximation result for $\operatorname{ReLU}$ activation was also given in [LLW21] with the following differences:
@@ 3 Proof of the Main Results
The objective of this section is to prove Theorems 2.7 and 2.9, following the three-step outline presented in Section 1.1. Our argument builds upon the methodology developed in [CLL21, LLW21, MLLR23, CLLZ24].
@@ 3.1.1 The Construction
    $u_{D}^{*}=\mathop{\operatorname{arg\,min}}_{u\in H_{0}^{1}(\Omega)}\mathcal{E}_{D}(u),\quad\mathcal{E}_{D}(u)\coloneqq\int_{\Omega}\left(\frac{1}{2}\nabla u\cdot A\nabla u+\frac{1}{2}cu^{2}-fu\right)\mathrm{d}x.$ (3.1)
@@
    $u_{N}^{*}=\mathop{\operatorname{arg\,min}}_{u\in H^{1}(\Omega)}\mathcal{E}_{N}(u),\quad\mathcal{E}_{N}(u)\coloneqq\int_{\Omega}\left(\frac{1}{2}\nabla u\cdot A\nabla u+\frac{1}{2}cu^{2}-fu\right)\mathrm{d}x.$ (3.2)
@@
The classical gradient descent method iteratively updates an initial guess by moving in the direction of the negative gradient of the objective function, aiming to converge to a minimizer. In the context of the PDEs we consider, Proposition 3.1 asserts that the weak solution is the unique minimizer of the associated energy functional. This naturally motivates the application of a gradient descent-type approach to solve the PDE. For the Dirichlet problem, we define the gradient of $\mathcal{E}_{D}$ as the map
@@
$\begin{split}\big(D_{H_{0}^{1}}\mathcal{E}_{D}(u),v\big)_{H^{1}(\Omega)}&=\lim_{t\to 0}\frac{\mathcal{E}_{D}(u+tv)-\mathcal{E}_{D}(u)}{t}\\ &=\int_{\Omega}\nabla u\cdot A\nabla v+cuv-fv\,\mathrm{d}x,\end{split}$ (3.3)
for all $u,v\in H_{0}^{1}(\Omega)$. Here we use the inner product on $H_{0}^{1}(\Omega)$ inherited from the standard inner product on $H^{1}(\Omega)$. The uniqueness of $D_{H_{0}^{1}}\mathcal{E}_{D}$ follows from the Riesz representation theorem and we call $D_{H_{0}^{1}}\mathcal{E}_{D}$ the Sobolev gradient of $\mathcal{E}_{D}$ in $H_{0}^{1}(\Omega)$. A similar argument applies to the energy functional $\mathcal{E}_{N}$ for the Neumann problem, and there exists a unique operator $D_{H^{1}}\mathcal{E}_{N}\colon H^{1}(\Omega)\to H^{1}(\Omega)$ such that
$\big(D_{H^{1}}\mathcal{E}_{N}(u),v\big)_{H^{1}(\Omega)}=\int_{\Omega}\nabla u\cdot A\nabla v+cuv-fv\,\mathrm{d}x,\quad\text{for all }u,v\in H^{1}(\Omega).$ (3.4)
We call $D_{H^{1}}\mathcal{E}_{N}$ the Sobolev gradient of $\mathcal{E}_{N}$ in $H^{1}(\Omega)$. We remark that the concept of the Sobolev gradient has been widely employed in the study of PDEs, including applications to the Gross–Pitaevskii eigenvalue problem [HP20, Zha22, CLLZ24].
Analogous to the fact that the gradient of a function vanishes at its local minimum, the Sobolev gradient of an energy functional should also vanish at its minimizer, i.e.,
@@ 3.1.2 Convergence Analysis
This subsection investigates the convergence rate of the Sobolev gradient flows to $u_{D}^{*}$ and $u_{N}^{*}$. In particular, we have the following theorem.
@@
Assume Assumption 2.1 holds and let $\alpha>0$ be sufficiently small. The Sobolev gradient flow (3.6) for the Dirichlet problem (1.3) converges exponentially in the $H^{1}(\Omega)$-norm to the weak solution $u_{D}^{*}$. More precisely, there exists a constant $0<\beta<1$ such that
@@
Proof.
We only present the proof for the Dirichlet problem (1.3), as the proof for the Neumann problem (1.4) follows analogously. Our goal is to establish that there exists a constant $0<\beta<1$ such that
@@
$\begin{split}&\|u_{t+1}-u_{D}^{*}\|_{H^{1}(\Omega)}^{2}=\|u_{t}-u_{D}^{*}-\alpha D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{H^{1}(\Omega)}^{2}\\ =&\|u_{t}-u_{D}^{*}\|_{H^{1}(\Omega)}^{2}+2\alpha\left(u_{D}^{*}-u_{t},D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\right)_{H^{1}(\Omega)}+\alpha^{2}\|D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{H^{1}(\Omega)}^{2}.\end{split}$ (3.8)
@@
Let $\lambda_{\max}\coloneqq\max\{a_{\max},c_{\max}\}$. Then, by combining (3.11), the definition of the Sobolev gradient in (3.3), and the Cauchy–Schwarz inequality, we obtain
$\|D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{H^{1}(\Omega)}^{2}$
$=$ $\left(D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t}),D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\right)_{H^{1}(\Omega)}-\left(D_{H_{0}^{1}}\mathcal{E}_{D}(u_{D}^{*}),D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\right)_{H^{1}(\Omega)}$
@@
$\leq$ $a_{\max}\|\nabla(u_{t}-u_{D}^{*})\|_{L^{2}(\Omega)}\|\nabla D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{L^{2}(\Omega)}$
$\qquad\qquad+c_{\max}\|u_{t}-u_{D}^{*}\|_{L^{2}(\Omega)}\|D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{L^{2}(\Omega)}$
$\leq$ $\lambda_{\max}\Big(\|\nabla(u_{t}-u_{D}^{*})\|_{L^{2}(\Omega)}\|\nabla D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{L^{2}(\Omega)}$
$\qquad\qquad+\|u_{t}-u_{D}^{*}\|_{L^{2}(\Omega)}\|D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{L^{2}(\Omega)}\Big)$
$\leq$ $\lambda_{\max}\left(\|\nabla(u_{t}-u_{D}^{*})\|_{L^{2}(\Omega)}^{2}+\|u_{t}-u_{D}^{*}\|_{L^{2}(\Omega)}^{2}\right)^{1/2}$
$\qquad\qquad\cdot\left(\|\nabla D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{L^{2}(\Omega)}^{2}+\|D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{L^{2}(\Omega)}^{2}\right)^{1/2}$
@@
$\beta\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\left(1-\lambda_{\min}\alpha+\lambda_{\max}^{2}\alpha^{2}\right)^{1/2}.$ (3.13)
Then for sufficiently small $\alpha>0$, we can ensure that $0<\beta<1$, which completes the proof. ∎
Let
and define the stepsize and contraction factor by
This choice of parameters is optimal in the sense that they minimize $\beta$ in (3.13).
@@
$\begin{split}\lambda_{\min}\|u_{D}^{*}\|_{H^{1}(\Omega)}^{2}&\leq\int_{\Omega}\nabla u_{D}^{*}\cdot A\nabla u_{D}^{*}\,\mathrm{d}x+\int_{\Omega}c|u_{D}^{*}|^{2}\,\mathrm{d}x\\ &=\int_{\Omega}fu_{D}^{*}\,\mathrm{d}x\leq\|f\|_{H^{-1}(\Omega)}\|u_{D}^{*}\|_{H^{1}(\Omega)}.\end{split}$ (3.15)
@@
then the iterates $u_{T}$ produced by both schemes (3.6) and (3.7) with stepsize $\alpha^{*}$ in (3.14) satisfy